OpenStudy (anonymous):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x).

OpenStudy (anonymous):

A. 7 + 4x B. 4 – x C. –x – 4 D. –4 – 7x

OpenStudy (anonymous):

f(x) + g(x) = (3x-2) + (6-4x) = -x + 4

OpenStudy (anonymous):

-x+4

OpenStudy (anonymous):

-x+4

OpenStudy (anonymous):

-x + 4 = 4 -x